# Non-locality in Quadrupolar Gravitational Radiation

### v1

@@ Abstract
General Relativity suffers for two main problems which have not yet been overcome: it predicts spacetime singularities and cannot be formulated as a perturbatively renormalizable theory. In particular, many attempts have been made for avoiding singularities, such as considering higher order or infinite derivative theories. The price to pay in both cases is to give up locality and therefore they are known altogether as non-local theories of gravity. In this paper, we investigate how to recognize the presence of non-local effects by exploiting the power emitted by gravitational waves in a binary system in presence of non-local corrections as $R\Box^{-1}R$ to the Hilbert-Einstein action. After solving the field equations in terms of the source stress-energy tensor $T_{\mu\nu}$ and obtaining the gravitational wave stress-energy pseudo-tensor, $\tau_{\mu\nu}$, we find that the General Relativity quadrupole formula is modified in a non-trivial way, making it feasible to find a possible signature of non-locality. Our final results on the gravitational wave stress-energy pseudo-tensor could also be applied to several astrophysical scenarios involving energy or momentum loss, potentially providing multiple tests for non-local deviations from General Relativity. We finally discuss the detectability of the massless transverse scalar mode, discovering that, although this radiation is extremely weak, in a small range around the model divergence, its amplitude could fall within the low-frequency Einstein Telescope sensitivity.
@@ 2 Non-local gravity as an extension of General Relativity
Defining $\psi_{0}\doteq 1+a\phi_{0}-\lambda_{0}$ and $\psi^{(1)}\doteq a\varphi-w$ its perturbation, and introducing the new field
@@
$K\Box w-\dfrac{6a^{2}}{\psi_{0}^{2}}\Box\varphi=0\ ,$ (2.17)
$K\Box\varphi-\dfrac{6a^{2}}{\psi_{0}^{2}}\Box w=0\ ,$ (2.18)
as already found in [43]. In Eqs. (2.17) and (2.18), we have defined the constant $K\doteq\dfrac{(6a-2\psi_{0})}{\psi_{0}^{2}}$. Notice that the last two are coupled differential equations for $\Box\varphi$ and $\Box w$. When $a=0$ they implies $\Box\varphi=0=\Box w$ and therefore $\varphi=0=w$, that is GR is recovered.
@@ 3.1 Explicit Solutions
$\theta_{\mu\nu}(t,\mathbf{x})=\dfrac{2G}{r}\dfrac{d^{2}Q_{ij}(t^{\prime})}{d^{2}t}\ ,$ (3.11)
@@ 3.2 The Emitted Power
where $w=\sqrt{\frac{GM}{4R^{3}}}$ is the keplerian angular frequency of the binary system. In general, for binary systems with different masses and on elliptic orbits of semi-major axis $R$ and eccentricity $e$, it turns out that
@@
where we have defined the function

### v2

@@ Abstract
General Relativity suffers for two main problems which have not yet been overcome: it predicts spacetime singularities and cannot be formulated as a perturbative renormalizable theory. In particular, many attempts have been made for avoiding singularities, such as considering higher order or infinite derivative theories. The price to pay in both cases is to give up locality and therefore they are known altogether as non-local theories of gravity. In this paper, we investigate how to recognize the presence of non-local effects by exploiting the power emitted by gravitational waves in a binary system in presence of non-local corrections as $R\Box^{-1}R$ to the Hilbert-Einstein action. After solving the field equations in terms of the source stress-energy tensor $T_{\mu\nu}$ and obtaining the gravitational wave stress-energy pseudo-tensor, $\tau_{\mu\nu}$, we find that the General Relativity quadrupole formula is modified in a non-trivial way, making it feasible to find a possible signature of non-locality. Our final results on the gravitational wave stress-energy pseudo-tensor could also be applied to several astrophysical scenarios involving energy or momentum loss, potentially providing multiple tests for non-local deviations from General Relativity. We finally discuss the detectability of the massless transverse scalar mode, discovering that, although this radiation is extremely weak, in a small range around the model divergence, its amplitude could fall within the low-frequency Einstein Telescope sensitivity.
@@ 2 Non-local gravity as an extension of General Relativity
An important issue is whether it is possible to eliminate or not the non-local terms from the Lagrangian (2.3) by redefining the metric tensor as done, for example, in Ref. [44]. This is not possible for the following mathematical and physical reasons. In fact, from a mathematical point of view, non-locality is linked to the integro-differential character of the field equations and therefore, from Eq.  (1.1), the non-local operators are integral operators where the value of a physical quantity at a point depends on its value assumed at other points modulated by the kernel. Thus it is not possible to eliminate the integral character of physical quantities through a redefinition of the metric which is a local tensor object. On the other hand, from a physical point of view, non-locality is in this theory an intrinsic property of gravity related to finite characteristic lengths and masses. Also if a "localization" process is performed (see e.g. [17, 30] ) auxiliary scalar fields emerge pointing out that non-locality introduces further degrees of freedom.
@@
Defining $\psi_{0}\doteq 1+a\phi_{0}-\lambda_{0}\neq 0$ and $\psi^{(1)}\doteq a\varphi-w$ its perturbation, and introducing the new field
@@
$K\Box w-\dfrac{6a}{\psi_{0}^{2}}\Box\varphi=0\ ,$ (2.17)
$K\Box\varphi-\dfrac{6a}{\psi_{0}^{2}}\Box w=0\ ,$ (2.18)
as already found in [43]. In Eqs. (2.17) and (2.18), we have defined the constant $K\doteq\dfrac{(6a-2\psi_{0})}{\psi_{0}^{2}}$. Notice that the last two are coupled differential equations for $\Box\varphi$ and $\Box w$. When $a=0$ they implies $\Box\varphi=0=\Box w$ and therefore $\varphi=0=w$, that is GR is recovered. The trivial solution of the system formed by the equations (2.17) and (2.18) is obtained by imposing that the determinant is different from zero, that is
$6a\neq\psi_{0},$ (2.19)
which implies
$\Box\varphi=\Box w=0,$ (2.20)
or
$\Box\psi^{(1)}=0\,.$ (2.21)
This means that $\phi^{(1)}$ is a massless scalar field added to the two tensor fields of the field Eq. (2.16). In other words, gravity linearly perturbed presents three degrees of freedom.
@@
As a final remark, we point out that combining relations in Eqs. (2.7), it turns out that
$\Box(\lambda+a\psi)=0,$ (2.26)
and then the first-order perturbation gives
$\Box(w+a\psi)=0,$ (2.27)
whose trivial solution is $w=-a\psi$. This means that the two fields $w$ and $\psi$ can be chosen as redefinitions of each other and that, therefore, it is necessary to add only one additional degree of freedom to the two standard tensor degrees of freedom of linearized GR. Based on this, the $f(R)$-like field $\psi^{(1)}$ can be written as $\psi^{(1)}=2a\varphi$. It is clear, from the above considerations and in the following sections, that the three scalar fields are related each other and non-local gravity implies only a further mode in GWs [36].
@@ 3.1 Explicit Solutions
$\theta_{ij}(t,\mathbf{x})=\dfrac{2G}{r}\dfrac{d^{2}Q_{ij}(t^{\prime})}{d^{2}t}\ ,$ (3.11)
@@ 3.2 The Emitted Power
where $w=\sqrt{\frac{GM}{4R^{3}}}$ is the keplerian angular frequency of the binary system. In general, for binary systems with different masses $m_{1}$ and $m_{2}$, on elliptic orbits of semi-major axis $R$ and eccentricity $e$, from the Peters - Mathews formula [49], it turns out that
@@
where $m=m_{1}+m_{2}$ is the total mass and $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ is the reduced mass, and we have defined the function